OpenStudy (anonymous):

PLSSS!!!! HELP!!!!! TOKENS!!!!! The half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass. Starting with 120 grams of a radioactive isotope, how much will be left after 3 half-lives?

OpenStudy (anonymous):

the half-life mean 1/2 of life so the 3half life will meaning 3*1/2 of life how you think it please ?

OpenStudy (anonymous):

yes, but wouldn't that give you \[ 120 \div \frac{ 3 }{ 2 }\] which gives you 80

OpenStudy (anonymous):

so this mean 120 halfed 3 time 120/2 = 60 60/2 = 30 30/2 = 15 so this will be understandably sure right easy now for you - yes ?

OpenStudy (anonymous):

Yeah thanks! but how would you write that in a function?